# A Syntax-Injected Approach for Faster and More Accurate Sentiment Analysis

### v1

@@ 3.1 Data
For polarity classification, we use a dataset of 1744 hotel reviews in English and 1438 reviews in Spanish from OpeNER which was created by Agerri et al. (2013). It contains an opinion mining corpus of hotel reviews for six languages (de, en, es, fr, it, nl) extracted from different booking sites from November 2012 to November 2013 Barnes et al. (2022). For our purposes, we only consider OpeNERes and OpeNERen. Each review is annotated with individual polarity expressions and their polarity (positive or negative) as demonstrated by a simple example of a review that only contains one polarity expression in Table 2 from Barnes et al. (2022). Since our problem of interest is polarity classification, we ignore the information of the respective holders and targets of polarity expressions, as well as the polarity strength such as “strong” and “standard” in the dataset. We also discard 350 reviews in English and 186 reviews in Spanish that do not contain any polarity expression at all, so our evaluation is conducted on the remaining reviews that do contain subjectivity. If the review is long and contains more than one polarity expression, e.g. “This hotel is expensive, but the staff is nice”, it is assigned a list of polarity values associated with each polarity expression such as [negative, positive] Barnes et al. (2022). Then, the aggregate polarity for the review as a whole is computed as the majority value in that list, or a third polarity (neutral) if there is no majority value (e.g. [negative, positive]). We thus have a task of predicting three polarity labels (neutral, positive, and negative) on this review dataset.
@@
In addition, in the case of Spanish, we also use data from the Rest-Mex 2023 sentiment dataset from TripAdvisor which contains 251.702 reviews in total, where Spanish-speaking Mexican Spanish tourists provide their evaluation or judgments on hotels, sightseeing, and restaurants (Álvarez Carmona et al., 2023). The Rest-Mex dataset contains a training dataset with polarity values accessible to the public, which was collected from various tourist destinations in Mexico, Cuba, and Colombia. The training data set includes labeled information about polarity, type of attraction, the country of origin for each opinion as well as the user-written title associated with the whole review, such as "recommendable!". This collection was obtained from tourists who shared their opinions on TripAdvisor between 2002 and 2023. Each opinion’s polarity is given on a five-star scale, i.e., as an integer between 1 and 5, where 1: Very bad, 2: Bad, 3: Neutral, 4: Good, 5: Very good. From the training dataset, we evaluate the polarity prediction task on a subset of 10 000 reviews, like the one shown in Table 1.222Note that our system has not been trained on these 10 000 reviews, as in fact it is an unsupervised rule-based system that needs no training at all. The reviews were not used in any way in the design of the rules or to tweak any system parameter. The reason for creating an evaluation dataset from the training set rather than evaluating on the test set is that the latter is not publicly available, so evaluating on it requires involving the Rest-Mex organizers. We do perform an evaluation on the test set for comparison with a Transformer-based system, as explained below, as that is a supervised system that is trained on the training set. We do not use titles for the polarity prediction task. In order to make the accuracy results from the polarity prediction task on the Rest-Mex reviews comparable with the accuracy results performed on the OpeNER datasets, as well as to facilitate straightforward and fair comparison with our VADER baseline, we convert the 5 polarity values associated with the 10 000 reviews from the Rest-Mex data set into 3 polarity labels (positive, negative, neutral) by merging the polarities 1 and 2 under the label "negative", polarities 4 and 5 under the label "positive" and polarity 3 under the label "neutral" (see Appendix B for conversion rules).
@@
For the accuracy comparison with Transformer-based models in the polarity prediction task, we also use the full test dataset of reviews which contains 107 863 reviews, since it would not be adequate to evaluate the Transformer-based supervised system on inputs on which it has been trained. The gold polarity values are not accessible to the public, which is why we send the polarity values output by the systems to the Rest-Mex organizers (Álvarez Carmona et al., 2023) who return to us the evaluation. In this polarity prediction task performed on the test dataset, we do not convert polarity values from 1 to 5 to three labels "positive", "negative" and "neutral", as we have no access to them, but directly evaluate accuracy on the five-star scale.
@@
Table 1: TripAdvisor Review in Spanish with corresponding polarity from the Rest-Mex dataset (Álvarez Carmona et al., 2023)
It is worth remarking that, except in the case of the Transformer-based model, we only use polarity prediction data for evaluation. With the aforementioned exception, the SA systems we are considering do not use a training set, as they are unsupervised with respect to polarity classification, not needing any training or undergoing any other form of tuning or tweaking with sentiment-annotated datasets. This kind of models have been shown to lag behind supervised models in accuracy when an in-domain corpus is available, but have more robust accuracy when it is not, a common case in practice Vilares et al. (2017).
@@
Table 2: Polarity expression and corresponding polarity from the English training dataset in (Barnes et al., 2021)
@@
It is important to note that the two datasets we are using have a significant length difference, although the topic of the data and domain are very similar as both datasets deal with reviews of touristic destinations such as hotels and restaurants. The RestMex dataset has longer reviews than the OpeNER dataset as measured by the mean of sentences per review as well as by the mean of tokens per sentence. The mean count of sentences per review in the Rest-Mex dataset: 3.75 vs. 1.06 in OpeNER. The mean count of tokens per sentence in the Rest-Mex: 23.77 vs. 16.38 in OpeNER.
@@
The datasets and code will be released on the GitHub repository [^3] https://github.com/chimran135/syntax-injected-sentiment-analysis. Moreover, the prototype  [^4] https://salsa.grupolys.org of the system is also available.
3.2 Methodology
Syntactic parsing
@@ 3.2 Methodology
To train the SELSP, we used DistilBERT-Base-Uncased as our sequence labeling classifier and evaluated the three sequence-labeling encodings mentioned in Section 2.4 on the validation sets of UD EWT English dataset (Silveira et al., 2014) and UD Spanish AnCora (Taulé et al., 2008) dataset. We chose the Relative Encoding (REL) as shown in Figure 1, as it obtained the best accuracy in these experiments.
@@
DistilBERT is a streamlined variant of BERT which maintains much of its performance while being smaller and faster. It requires less computational resources and suits well for the tasks where efficiency and speed are crucial, such as dependency parsing. We fine-tuned the DistilBERT model (Sanh et al., 2019) from the pretrained checkpoint (DistilBERT-base-uncased) available in the HuggingFace library using UD-EWT English dataset and UD Spanish AnCora dataset. DistilBERT consists of a pre-trained transformer encoder, and during fine-tuning, only the parameters of this encoder are updated. The basic idea behind fine-tuning was to leverage the knowledge learned by DistilBERT during pre-training on a large corpus and to adapt it for the sequence labeling task.
@@
To fine-tune the SELSP, the original training, validation and test data splits from the considered English and Spanish UD datasets were first converted into sequence labels using the CoDeLin framework (Section 2.4). Then, the model was trained for 30 epochs on the resulting training set, and validation was performed to adjust the learned weights and biases after each epoch. Table  3 describes the other hyperparameters used to train the model. The same hyperparameters were used to train both English and Spanish models.
@@
The training process results into a model that, given an input sentence, predicts the labels encoding the dependency heads and relations in our chosen encoding, in this case the Relative Encoding (REL). We then use the CoDeLin sequence labeling framework to decode these rules produced by the SELSP into the CoNLL-U format that is standard for Universal Dependencies. With this, the accuracy of the parse trees produced by the fine-tuned SELSP was evaluated using the CoNLL 2018 standard evaluation script (Zeman et al., 2018).
@@
The CoNLL-UD format represents a dependency graph in a particular format according to which tokens are annotated linguistically concerning their lemma, POS, morphology, the location of their syntactic head expressed as an absolute token index, and the label of the dependency relation such as subject relation (nsubj), object relation (obj), determiner (det), adverbial modification (advmod), conjunction (cc), conjoined element (conj), complementizer (xcomp), or punctuation (punct), as shown in Figure 2. In this format, word tokens are indexed with integers starting from 1, and all the information about a given token is in a separate line. The information of each dependency relation is included in the line associated with the dependent, by specifying the index of its head together with the relation type. For instance, in the example in the figure, the head of the word token with the integer 2 is the word 3 (the main verb), with relation type “nsubj” indicating that “I” is the subject of “found”; and the modifier “very” at position 10 is modifying the word "kind" at position 11. The sentence’s syntactic root (the main verb "found" in the example) has no head, so its associated head index is set to zero. This information generated by the dependency parser is used to identify polarity shifters: for example, since “very” is a modifier of “kind”, it will increase the positive polarity associated with “kind”. The polarity prediction demonstrated between the dependency graph and the CoNLL-UD format in Figure 2 will be explained below.
@@
Figure 2: Polarity Prediction based on a UD-based dependency tree from (Kellert et al., 2023)
@@
Table 3: Fine-Tuning the SELSP
Polarity classification and sentiment dictionaries
We then performed the integration of the SELSP with the sentiment analysis code with the polarity rules of (Kellert et al., 2023). For this purpose, we also developed a mapping function to convert the CoNLL-U format to the Stanza output format, so that the same sentiment analysis code (initially designed for Stanza) may also run for SELSP (or other parsers producing CoNLL-U output) interchangeably.
@@
The original sentiment analysis system of (Kellert et al., 2023) contains information on how to calculate polarity scores of reviews based on sentiment words from the sentiment dictionary SO-CAL (Brooke et al., 2009) and UD-based syntactic rules for Spanish, using syntactic parse trees obtained from Stanza (Kellert et al., 2023). Apart from incorporating the option of parsing with SELSP as described above, we also extended the code to handle both English and Spanish, as well as to be able to incorporate different sentiment dictionaries apart from SO-CAL. We now describe the extensions made.
@@
We use a set of English negation words and intensifiers like "very" or weakeners like "few" to process English sentiment data, which can be found in § Appendix A. The set of negation words we use for processing Spanish data is integrated into the code from (Kellert et al., 2023).
@@
We use several English and Spanish sentiment dictionaries. For the English sentiment dataset, we use the following dictionaries: 1) SO-CAL, which is available in Spanish and English (Brooke et al., 2009), 2) VADER Lexicon (VADERdict), which is only available in English (Hutto and Gilbert, 2014), 3) an aggregate version that we obtain by adding to SO-CAL the sentiment words from VADERdict that are not initially covered by SO-CAL, and 4) the English dictionary from BabelSenticNet Vilares et al. (2018). For the two Spanish sentiment datasets (OpeNER reviews and the 10 000 Spanish TripAdvisor reviews from RestMex), we use 1) the Spanish version of SO-CAL, 2) an aggregate version that combines SO-CAL with words from VADERdict that are missing in the English SO-CAL that we translate into Spanish via DeepL (SO-CAL + sp. VADERdict), and two automatically-extracted datasets: 3) words with low ambiguity extracted from titles from the Rest-Mex training dataset Kellert et al. (2024) and 4) an aggregate version that combines SO-CAL with words from VADERdict and Rest-Mex dictionary.
@@
We now provide some details on how these two latter dictionaries have been generated. We measure the polarity ambiguity of words by the standard deviation of polarity judgments (sdv) Kellert et al. (2024). The authors of said paper motivate their method of constructing sentiment dictionaries by using words with low sdv from titles as follows. Titles of reviews represent short summaries of long reviews and the words from these summaries often function like sentiment words, e.g., "Recommendable!", "Excellent!", etc. However, sometimes titles also represent topics of reviews, like "Last holidays", and the words from these topic-titles cannot be considered as sentiment words. In this case, as the authors show (Kellert et al., 2024), words from these titles tend to be associated with a high sdv, since people will provide different, sometimes opposite judgments, to reviews containing that kind of topic words. For this reason, the authors consider only words with low sdv as good candidates for building sentiment words out of titles. Thus, we only select words with lower polarity variation measured by sdv (<0.6). For the experiments in this study, we use two different word lists. One word list contains words with low sdv from short titles (ShortT), which are those that contain only one word such as “Excellent!”, and the other list contains words with low sdv from short and long titles (ShortT+LongT), where long titles are those with multiple words. It is important to note that the accuracy calculation of the polarity prediction task based on the Rest-Mex dataset was performed on whole review texts excluding their titles. As a consequence, the sentiment dictionary created from the titles of the Rest-Mex dataset does not influence the accuracy of the polarity prediction task based on reviews from the Rest-Mex dataset.
@@
Regarding aggregation of dictionaries (like when we combine SO-CAL with VADERdict), the issue can arise that two dictionaries have different polarities for the same word. Throughout the paper, we will consider that SO-CAL is the basic dictionary to which we add words with corresponding polarities from other dictionaries that are not contained in SO-CAL. Thus, if SO-CAL has a lemma/word with a given polarity, but we combine it with another dictionary (such as VADERdict or the dictionaries extracted from Rest-Mex data) where the same lemma/word has a different polarity, we always keep the score assigned by SO-CAL.
@@
For all these sentiment dictionaries, we compare the polarity results produced by our system that combines SELSP and syntactic rules with two main baselines. The first one is the version of the same SA rules that used Stanza-based syntactic parsing as their backbone instead (Kellert et al., 2023). The second main baseline is the shallow heuristics-based approach VADER, which is a widely used in practical applications of SA to research and industry use cases (see e.g. Lasser et al. (2023); Deas et al. (2023); Cookson et al. (2023) for some recent examples).
@@
We also compare our results from the Spanish polarity prediction task with the state-of-the-art Transformer-based model from the javier-alonso team that achieved the second rank in the polarity prediction task in the Rest-Mex shared task 2023 (Álvarez Carmona et al., 2023; Alonso-Mencía, 2023). The javier-alonso team (Alonso-Mencía, 2023) used two variants of the RoBERTa model, namely roberta-base-bne and twitter-xlm-robertabase. They balanced the effect of minority classes (such as polarity 1 and 2) through oversampling strategies. The reason why we used the Transformer-based model from the second highest ranked team and not the first is that their source code was easily accessible to us. Note, however, that the results from both teams in the shared task were very similar, with a difference of only one percentage point in the polarity classification task. Thus, we can assume that the results of this system are a reasonable representation of the state of the art.
@@
For the comparison with the Transformer-based model, we use the following dictionaries: 1) words with low ambiguity extracted from short and long titles from the Rest-Mex training data as explained before Kellert et al. (2024) and 2) a combination of dictionaries: the Spanish version of SO-CAL, together with words from the VADER dictionary that are missing in English SO-CAL that we translate into Spanish via DeepL (sp. VADERdict), and words with low ambiguity extracted from short and long titles from the Rest-Mex training data (Kellert et al., 2024).
@@
For the evaluation of the systems on the polarity prediction task, we converted the outputs of these systems into class labels for ternary classification ([positive, neutral, negative]) (see § Appendix B on conversion rules), which we then compared with the gold polarity classes obtained from the datasets in the same format (Section 3.1). As an exception, for the comparison with the Transformer-based model, we perform the evaluation directly on the five-star scale (i.e., 5-class classification) due to the restrictions inherent to this evaluation: as mentioned before, the model has been trained and adjusted on the training set, so we must use the test set to which we do not have direct access, performing instead the evaluation by sending system outputs to the task organizers. Therefore, no conversion is possible.
3.3 Results
We first report our results for dependency parsing, which serves as the basis for the subsequent polarity classification. The parsing results in Table 4 show that our SELSP achieves over 75% accuracy (in terms of LAS, i.e., percentage of words for which both the head and the dependent are correctly predicted) in syntactic parsing on the UD EWT English and UD Spanish AnCora treebanks. While this is lower than state-of-the-art accuracy for these treebanks, it is worth reminding that syntax-based polarity classification does not need high parsing accuracy (Gómez-Rodríguez et al., 2017) so our goal here is to optimize for speed (hence using a sequence-labeling approach and a small distilled language model, decisions that sacrifice accuracy in favor of efficiency).
@@ 3.3 Results
Table  6, Table  7, and Table  8 show the accuracy results for SELSP, Stanza, and VADER for different datasets and different dictionaries (see Section 3.2). Compared to the baselines, the results show that SELSP achieves clearly better results than VADER across all language, dataset and dictionary combinations. When comparing SELSP to Stanza, we observe that the accuracy of the system using the SELSP parser is not only on par with the version using Stanza, but even outperforms it in most cases (9 dataset-dictionary combinations where SELSP provides better accuracy than Stanza, vs. 3 where Stanza is better). This is particularly interesting, given that our goal in replacing Stanza with SELSP was to optimize for efficiency, and the SELSP parser prioritizes speed over accuracy (Table 4) and in particular less accurate than Stanza. The fact that this does not result on any loss of accuracy on the end task is a confirmation of the previous observation by Gómez-Rodríguez et al. (2017) that syntax-based polarity classification performance no longer improves with parsing accuracy once a “good enough” parsing accuracy threshold is achieved. Clearly, the accuracies provided by SELSP in Table 4 are above that threshold for our datasets under consideration, so our results confirm that it is clearly beneficial to use parsing as sequence labeling in this scenario.
@@
Table 6: Comparison between SELSP, Stanza and VADER in accuracy (percent) in the polarity prediction task performed on the English dataset from OpeNER across several dictionaries (see Section 3.2) based on three polarity labels "neutral", "positive" and "negative"
@@
Table 7: Comparison between SELSP, Stanza and VADER in accuracy (percent) in the polarity prediction task performed on the Spanish dataset from OpeNER across several dictionaries based on three polarity labels "neutral", "positive" and "negative"

### v2

@@ Introduction
For polarity classification, we use a dataset of 1743 hotel reviews in English and 1438 reviews in Spanish from OpeNER [^4] https://github.com/jerbarnes/semeval22_structured_sentiment which was created by Agerri et al., (2013). It contains an opinion mining corpus of hotel reviews for six languages (de, en, es, fr, it, nl) extracted from different booking sites from November 2012 to November 2013 (Barnes et al.,, 2022). For our purposes, we only consider the English and Spanish datasets, OpeNERes and OpeNERen. Each review is annotated with individual polarity expressions and their polarity (positive or negative) as demonstrated by a simple example of a review that only contains one polarity expression in Table 2 from (Barnes et al.,, 2022). Since our problem of interest is polarity classification, we ignore the information of the respective holders and targets of polarity expressions, as well as the polarity strength such as “strong” and “standard” in the dataset. We also discard 350 reviews in English and 186 reviews in Spanish that do not contain any polarity expression at all (no annotation provided in the datasets), so our evaluation is conducted on the remaining reviews that do contain subjectivity. If the review is long and contains more than one polarity expression, e.g. “This hotel is expensive, but the staff is nice”, it is assigned a list of polarity values associated with each polarity expression such as [negative, positive] (Barnes et al.,, 2022). Then, the aggregate polarity for the review as a whole is computed as the majority value in that list, or a third polarity (neutral) if there is no majority value (e.g. [negative, positive]). We thus have a task of predicting three polarity labels (neutral, positive, and negative) on this review dataset.
@@
In addition, in the case of Spanish, we also use data from the Rest-Mex 2023 sentiment dataset [^5] https://sites.google.com/cimat.mx/rest-mex2023 from TripAdvisor which contains 251 702 reviews in total, where Spanish-speaking Mexican Spanish tourists provide their evaluation or judgments on hotels, sightseeing, and restaurants (Álvarez-Carmona et al.,, 2023). The Rest-Mex dataset contains a training dataset with polarity values accessible to the public, which was collected from various tourist destinations in Mexico, Cuba, and Colombia. The training data set includes labeled information about polarity, type of attraction, the country of origin for each opinion as well as the user-written title associated with the whole review, such as “recommendable!”. This collection was obtained from tourists who shared their opinions on TripAdvisor between 2002 and 2023. Each opinion’s polarity is given on a five-star scale, i.e., as an integer between 1 and 5, where 1: Very bad, 2: Bad, 3: Neutral, 4: Good, 5: Very good. From the training dataset, we evaluate the polarity prediction task on a subset of 10 000 reviews, like the one shown in Table 1.666Note that our system has not been trained on these 10 000 reviews, as in fact it is an unsupervised rule-based system that needs no training at all. The reviews were not used in any way in the design of the rules or to tweak any system parameter. The reason for creating an evaluation dataset from the training set rather than evaluating on the test set is that the latter is not publicly available, so evaluating on it requires involving the Rest-Mex organizers. We do perform an evaluation on the test set for comparison with a Transformer-based system, as explained below, as that is a supervised system that is trained on the training set. We do not use titles for the polarity prediction task. In order to make the accuracy results from the polarity prediction task on the Rest-Mex reviews comparable with the accuracy results performed on the OpeNER datasets, as well as to facilitate straightforward and fair comparison with our VADER baseline, we convert the 5 polarity values associated with the 10 000 reviews from the Rest-Mex data set into 3 polarity labels (positive, negative, neutral) by merging the polarities 1 and 2 under the label “negative”, polarities 4 and 5 under the label “positive” and polarity 3 under the label “neutral” (see Appendix B for conversion rules).
@@
For the accuracy comparison with Transformer-based models in the polarity prediction task, we focus on the Rest-Mex dataset where we use the full test set containing 107 863 reviews, since it would not be adequate to evaluate the Transformer-based supervised system on inputs on which it has been trained. The gold polarity values are not accessible to the public, which is why we send the polarity values output by the systems to the Rest-Mex organizers (Álvarez-Carmona et al.,, 2023) who return the evaluation to us. In this polarity prediction task performed on the test dataset, we do not convert polarity values from 1 to 5 to three labels ”positive”, ”negative” and ”neutral”, as we have no access to them, but directly evaluate accuracy on the five-star scale.
@@
Table 1: TripAdvisor Review in Spanish with corresponding polarity from the Rest-Mex dataset (Álvarez-Carmona et al.,, 2023)
It is worth remarking that, except in the case of the Transformer-based model, polarity data from the sentiment datasets is only used for evaluation. With the aforementioned exception, the SA systems we are considering do not use a training set, as they are unsupervised with respect to polarity classification, not needing any training or undergoing any other form of tuning or tweaking with sentiment-annotated datasets. This kind of models have been shown to lag behind supervised models in accuracy when an in-domain corpus is available, but have more robust accuracy when it is not, a common case in practice (Vilares et al.,, 2017).
@@
Table 2: Polarity expression and corresponding polarity from the English training dataset in (Barnes et al.,, 2021)
@@
It is important to note that the two datasets we are using have a significant text length difference, although the topic of the data and domain are very similar as both datasets deal with reviews of touristic destinations such as hotels and restaurants. The RestMex dataset has longer reviews than the OpeNER dataset as measured by the mean of sentences per review as well as by the mean of tokens per sentence. The mean count of sentences per review in the Rest-Mex dataset is 3.75, vs. 1.06 in OpeNER. The mean count of tokens per sentence in the Rest-Mex dataset is 23.77, vs. 16.38 in OpeNER.
@@
The datasets and code are available in the Zenodo repository [^7] https://doi.org/10.5281/zenodo.15323755. Moreover, the prototype  [^8] https://salsa.grupolys.org of the system is also available.
Methodology
The sentiment analysis methodology consists of two modules; a sequence labeling-based dependency parser (SELSP) to generate dependency trees followed by a sentiment analysis engine which uses sentiment dictionaries and syntactic rules for sentiment polarity classification from the said dependency trees.
Syntactic parsing (SELSP Parser)
@@
To train the SELSP, we used DistilBERT-Base-Uncased as our sequence labeling classifier and evaluated the three sequence-labeling encodings mentioned in Section “Dependency parsing approaches” on the validation sets of the UD EWT English dataset (Silveira et al.,, 2014) and the UD Spanish AnCora (Taulé et al.,, 2008) dataset. We chose the Relative Encoding (REL) as shown in Figure 1, as it obtained the best accuracy in these experiments.
@@
DistilBERT is a streamlined variant of BERT which maintains much of its performance while being smaller and faster. It requires less computational resources and suits well for the tasks where efficiency and speed are crucial, such as dependency parsing. We fine-tuned the DistilBERT model (Sanh et al.,, 2019) from the pretrained checkpoint (DistilBERT-base-uncased) available in the HuggingFace library using the UD-EWT English dataset and the UD Spanish AnCora dataset. DistilBERT consists of a pre-trained transformer encoder, and during fine-tuning, only the parameters of this encoder are updated. The basic idea behind fine-tuning was to leverage the knowledge learned by DistilBERT during pre-training on a large corpus and to adapt it for the sequence labeling task.
@@
To fine-tune the SELSP, the original training, validation and test data splits from the considered English and Spanish UD datasets were first converted into sequence labels using the CoDeLin framework (Section “Dependency parsing approaches”). Then, the model was trained for 30 epochs on the resulting training set, and validation was performed to adjust the learned weights and biases after each epoch. Table  3 describes the other hyperparameters used to train the model. The same hyperparameters were used to train both English and Spanish models. The pipeline of SELSP parser is given in Figure 2.
Figure 2: The pipeline of SELSP parser
@@
The training process results into a model that, given an input sentence, predicts the labels encoding the dependency heads and relations in our chosen encoding, in this case the Relative Encoding (REL). We then use the CoDeLin sequence labeling framework to decode these rules produced by the SELSP into the CoNLL-U format that is standard for Universal Dependencies. With this, the accuracy of the parse trees produced by the fine-tuned SELSP was evaluated using the CoNLL 2018 standard evaluation script (Zeman et al.,, 2018).
@@
The CoNLL-UD format represents a dependency graph in a particular format according to which tokens are annotated linguistically concerning their lemma, POS, morphology, the location of their syntactic head expressed as an absolute token index, and the label of the dependency relation such as subject relation (nsubj), object relation (obj), determiner (det), adverbial modification (advmod), conjunction (cc), conjoined element (conj), complementizer (xcomp), or punctuation (punct), as shown in Figure 3. In this format, word tokens are indexed with integers starting from 1, and all the information about a given token is in a separate line. The information of each dependency relation is included in the line associated with the dependent, by specifying the index of its head together with the relation type. For instance, in the example in the figure, the head of the word token with the integer 2 is the word 3 (the main verb), with relation type “nsubj” indicating that “I” is the subject of “found”; and the modifier “very” at position 10 is modifying the word “kind” at position 11. The sentence’s syntactic root (the main verb “found” in the example) has no head, so its associated head index is set to zero. This information generated by the dependency parser is used to identify polarity shifters: for example, since “very” is a modifier of “kind”, it will increase the positive polarity associated with “kind”. The polarity prediction demonstrated between the dependency graph and the CoNLL-UD format in Figure 3 will be explained below.
@@
Figure 3: Polarity Prediction based on a UD-based dependency tree from (Kellert et al.,, 2023)
@@
Table 3: Hyperparameters to fine-tune the SELSP
Sentiment Analysis Engine for Polarity Classification
Once the SELSP is trained, its output (dependency tree) is fed into the sentiment analysis engine which uses syntactic rules  (Kellert et al.,, 2023) to calculate sentiment polarity from said dependency trees. The overall architecture of the sentiment analysis system is provided in Figure 4.
Figure 4: Overall architecture of the sentiment analysis system
@@
The original sentiment analysis system of (Kellert et al.,, 2023) contains information on how to calculate polarity scores of reviews based on sentiment words from the sentiment dictionary SO-CAL (Brooke et al.,, 2009) and UD-based syntactic rules for Spanish, using syntactic parse trees obtained from Stanza (Kellert et al.,, 2023). Apart from incorporating the option of parsing with SELSP as described above, we also extended the system to handle both English and Spanish, as well as to be able to incorporate different sentiment dictionaries apart from SO-CAL. We now describe the extensions made.
@@
We use a set of English negation words and intensifiers like “very” or weakeners like “few” to process English sentiment data, which can be found in § Appendix A. The set of negation words we use for processing Spanish data is integrated into the code from (Kellert et al.,, 2023).
@@
We use several English and Spanish sentiment dictionaries. For the English sentiment dataset, we use the following dictionaries: 1) SO-CAL, which is available in Spanish and English (Brooke et al.,, 2009), 2) VADER Lexicon (VADERdict), which is only available in English (Hutto and Gilbert,, 2014), 3) an aggregate version that we obtain by adding to SO-CAL the sentiment words from VADERdict that are not initially covered by SO-CAL, and 4) the English dictionary from BabelSenticNet (Vilares et al.,, 2018). For the two Spanish sentiment datasets (OpeNER reviews and the 10 000 Spanish TripAdvisor reviews from RestMex), we use 1) the Spanish version of SO-CAL, 2) an aggregate version that combines SO-CAL with words from VADERdict that are missing in the English SO-CAL that we translate into Spanish via DeepL (SO-CAL + sp. VADERdict), and two automatically-extracted datasets: 3) words with low ambiguity extracted from titles from the Rest-Mex training dataset (Kellert et al.,, 2024) and 4) an aggregate version that combines SO-CAL with words from VADERdict and Rest-Mex dictionary.
@@
We now provide some details on how these two latter dictionaries have been generated. We measure the polarity ambiguity of words by the standard deviation of polarity judgments (sdv) (Kellert et al.,, 2024). The authors of said paper motivate their method of constructing sentiment dictionaries by using words with low sdv from titles as follows. Titles of reviews represent short summaries of long reviews and the words from these summaries often function like sentiment words, e.g., “Recommendable!”, “Excellent!”, etc. However, sometimes titles also represent topics of reviews, like “Last holidays”, and the words from these topic-titles cannot be considered as sentiment words. In this case, as the authors show (Kellert et al.,, 2024), words from these titles tend to be associated with a high sdv, since people will provide different, sometimes opposite judgments, to reviews containing that kind of topic words. For this reason, the authors consider only words with low sdv as good candidates for building sentiment words out of titles. Thus, we only select words with lower polarity variation measured by sdv ($<0.6$). For the experiments in this study, we use two different word lists. One word list contains words with low sdv from short titles (ShortT), which are those that contain only one word such as “Excellent!”, and the other list contains words with low sdv from short and long titles (ShortT+LongT), where long titles are those with multiple words. It is important to note that the accuracy calculation of the polarity prediction task based on the Rest-Mex dataset was performed on whole review texts excluding their titles. As a consequence, the sentiment dictionary created from the titles of the Rest-Mex dataset does not influence the accuracy of the polarity prediction task based on reviews from the Rest-Mex dataset.
@@
Regarding aggregation of sentiment dictionaries (like when we combine SO-CAL with VADERdict), the issue can arise that two dictionaries have different polarities for the same word. Throughout the paper, we will consider that SO-CAL is the basic dictionary to which we add words with corresponding polarities from other dictionaries that are not contained in SO-CAL. Thus, if SO-CAL has a lemma/word with a given polarity, but we combine it with another dictionary (such as VADERdict or the dictionaries extracted from Rest-Mex data) where the same lemma/word has a different polarity, we always keep the score assigned by SO-CAL.
@@
For all these sentiment dictionaries, we compare the polarity results produced by our system that combines SELSP and sentiment analysis engine with two main baselines. The first one is the version of the same SA rules that used Stanza-based syntactic parsing as their backbone instead (Kellert et al.,, 2023). The second main baseline is the shallow heuristics-based approach VADER, which is a widely used in practical applications of SA to research and industry use cases (see e.g. (Lasser et al.,, 2023; Deas et al.,, 2023; Yuliansyah et al.,, 2024; Barik and Misra,, 2024; Lahande et al.,, 2025) for some recent examples).
@@
We also compare our results from the Spanish polarity prediction task with the state-of-the-art Transformer-based model from the javier-alonso team that achieved the second rank in the polarity prediction task in the Rest-Mex shared task 2023 (Álvarez-Carmona et al.,, 2023; Alonso-Mencía,, 2023). The javier-alonso team (Alonso-Mencía,, 2023) used two variants of the RoBERTa model, namely roberta-base-bne and twitter-xlm-roberta-base. They balanced the effect of minority classes (such as polarity 1 and 2) through oversampling strategies. The reason why we used the Transformer-based model from the second highest ranked team and not the first is that their source code was easily accessible to us. Note, however, that the results from both teams in the shared task were very similar, with a difference of only one percentage point in the polarity classification task. Thus, we can assume that the results of this system are a reasonable representation of the state of the art.
@@
For the comparison with the Transformer-based model, we use the following dictionaries: 1) words with low ambiguity extracted from short and long titles from the Rest-Mex training data as explained before (Kellert et al.,, 2024) and 2) a combination of dictionaries: the Spanish version of SO-CAL, together with words from the VADER dictionary that are missing in English SO-CAL that we translate into Spanish via DeepL (sp. VADERdict), and words with low ambiguity extracted from short and long titles from the Rest-Mex training data (Kellert et al.,, 2024).
@@
For the evaluation of the systems on the polarity prediction task, we converted the outputs of these systems into class labels for ternary classification ([positive, neutral, negative]) (see § Appendix B on conversion rules), which we then compared with the gold polarity classes obtained from the datasets in the same format (Section “Data”). As an exception, for the comparison with the Transformer-based model, we perform the evaluation directly on the five-star scale (i.e., 5-class classification) due to the restrictions inherent to this evaluation: as mentioned before, the model has been trained and adjusted on the training set, so we must use the test set to which we do not have direct access, performing instead the evaluation by sending system outputs to the task organizers. Therefore, no conversion is possible.
Results and Discussion
We first report our results for dependency parsing, which serves as the basis for the subsequent polarity classification. The parsing results in Table 4 show that our SELSP achieves over 75% accuracy (in terms of LAS, i.e., percentage of words for which both the head and the dependent are correctly predicted) in syntactic parsing on the UD EWT English and UD Spanish AnCora treebanks. While this is lower than state-of-the-art accuracy for these treebanks, it is worth reminding that syntax-based polarity classification does not need high parsing accuracy (Gómez-Rodríguez et al.,, 2019) so our goal here is to optimize for speed (hence using a sequence-labeling approach and a small distilled language model, decisions that sacrifice accuracy in favor of efficiency).
@@
Table 6, Table 7, and Table 8 show the accuracy results for SELSP, Stanza, and VADER for different datasets and different dictionaries (see Section “Methodology”). Compared to the baselines, the results show that SELSP achieves clearly better results than VADER across all language, dataset and dictionary combinations. When comparing SELSP to Stanza, we observe that the accuracy of the system using the SELSP parser is not only on par with the version using Stanza, but even outperforms it in most cases (9 dataset-dictionary combinations where SELSP provides better accuracy than Stanza, vs. 3 where Stanza is better). This is particularly interesting, given that our goal in replacing Stanza with SELSP was to optimize for efficiency, and the SELSP parser prioritizes speed over accuracy (Table 4) and in particular is less accurate than Stanza. The fact that this does not result on any loss of accuracy on the end task is a confirmation of the previous observation by Gómez-Rodríguez et al., (2019) that syntax-based polarity classification performance no longer improves with parsing accuracy once a “good enough” parsing accuracy threshold is achieved. Clearly, the accuracies provided by SELSP in Table 4 are above that threshold for our datasets under consideration, so our results confirm that it is clearly beneficial to use parsing as sequence labeling in this scenario.
@@
Table 6: Comparison between SELSP, Stanza and VADER in accuracy (percent) in the polarity prediction task performed on the English dataset from OpeNER across several dictionaries (see Section “Methodology”) based on three polarity labels ”neutral”, ”positive” and ”negative”
@@
Table 7: Comparison between SELSP, Stanza and VADER in accuracy (percent) in the polarity prediction task performed on the Spanish dataset from OpeNER across several dictionaries based on three polarity labels ”neutral”, ”positive” and ”negative”